Solve


solve (3x^2+2xy^2)dx+(2x^2y)dy=0 where y(2)=-3


1
Expert's answer
2021-06-29T07:50:11-0400
"(3x^2+2xy^2)dx+(2x^2y)dy=0"

"P=3x^2+2xy^2, \dfrac{\partial P}{\partial y}=4xy"

"Q=2x^2y, \dfrac{\partial Q}{\partial x}=4xy"

"\dfrac{\partial P}{\partial y}=4xy=\dfrac{\partial Q}{\partial x}"

"\dfrac{\partial u}{\partial x}=P(x, y), \dfrac{\partial u}{\partial y}=Q(x,y)"

"u(x,y)=\int(3x^2+2xy^2)dx+\varphi(y)"

"=x^3+x^2y^2+\varphi(y)"

"\dfrac{\partial u}{\partial y}=2x^2y+\varphi'(y)=2x^2y"

"\varphi'(y)=0"

"\varphi(y)=C_1"

The general solution of the differential equation


"(3x^2+2xy^2)dx+(2x^2y)dy=0"

is given by


"x^3+x^2y^2=C"

 y(2)=-3


"(2)^3+(2)^2(-3)^2=C"

"C=44"

The solution of the IVP is


"x^3+x^2y^2=44"